OpenStudy (anonymous):

x^2-2x-35=0 use the quadratic formula to solve the equation

OpenStudy (lgbasallote):

...so use the quadratic formula...what's the problem?

OpenStudy (anonymous):

This is in the form \(y = ax^2 + bx + c\). THis is the quadratic formula: \[x = \frac{-b ± \sqrt{b^2 - 4ac}}{2a}\]Plug in a, b, and c based on the first thing I said.

OpenStudy (anonymous):

i dont know how to solve

OpenStudy (anonymous):

Literally, just plug in numbers and simplify... \[x = \frac{2 ± \sqrt{(-2)^2 - 4(1)(-35)}}{2(1)} \implies?\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

10/2

OpenStudy (anonymous):

THat's one. What about the other? \[x = \frac{2 ± 12}{2} \implies x = 1 ± 6\]

OpenStudy (anonymous):

So what are your two answers now?

OpenStudy (anonymous):

sorry